VT (9 pts) Suppose the random variable Y has a geometric distribution with p Find the probability distribution of V arameter p. Let V- 4.
Answer #1

The PMF of Y with parameter p . Pleft ( Y=y ight )=left ( 1-p ight )^{y-1}p;y=1,2,3,4,..., Since Y has geometric distribution with parameter p. Now consider the tranformation V=rac{sqrt[3]{Y}}{3} . Therefore, the PMF of the tranformatiom, 3 Y3 is

P (V-u) = (1-p)27U 3-1 p: 271,3-1. 2. 3. 4.